OpenStudy (anonymous):

if cos x=2/3 and x is in quadrant 4 then sin x /2= - 1/3 1/3 -sqt1/6 sqt1/6

OpenStudy (ikram002p):

u mean \(\large \sin (\frac{x}{2})\) or \(\Huge \frac{\sin x}{2}\) ??

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

Does anyone know how to solve this :S

OpenStudy (anonymous):

have you tried calculating what alpha is equal to, putting it into the cast diagram and finding out theta?

OpenStudy (anonymous):

I don't know how to do that :/

OpenStudy (anonymous):

on your calculator calculate the inverse sin of x and use the cast diagram to calculate what theta is equal to by cycling round the cast diagram

OpenStudy (anonymous):

√(1/6)